# Electromagnetic moments in the Sn-Gd region determined within the nuclear DFT

### v1

@@ Abstract
Within the nuclear DFT framework, employing the Skyrme UNEDF1 functional and incorporating pairing correlations, we determined the spectroscopic electric quadrupole and magnetic dipole moments of the $\nu 11/2^{-}$ and $\pi 7/2^{+}$ configurations in heavy, deformed, open-shell odd nuclei with $50\leq Z\leq 64$. The notions of self-consistent shape and spin polarisations due to odd nucleons, which are responsible for generating total electric quadrupole and magnetic dipole moments, were transformed into detailed computational procedures. The alignment of intrinsic angular momentum along the axis of axial symmetry, necessitating signature and time-reversal symmetry breaking, followed by the restoration of rotational symmetry, proved to be essential components of the method. In contrast, the restoration of particle number symmetry yields modifications of only about 1%. With the isovector spin-spin terms of the functional previously adjusted in near doubly magic nuclei across the mass chart, the calculations were parameter-free. Effective charges and $g$-factors were not employed. A reasonably good agreement was achieved between the calculated and measured electric quadrupole moments. A similarly fair description of the magnetic dipole moments was obtained for the intruder configurations $\nu 11/2^{-}$ alongside a poor description of those for $\pi 7/2^{+}$.
@@ 1 Introduction
In the present work, we expand the scope of the results presented in [5] to encompass the region of lighter elements ranging from tin to gadolinium. Within nuclear density functional theory (DFT), we determine the spectroscopic electric quadrupole and magnetic dipole moments of the nuclei depicted in figure 1. Our dual focus here is on studying (i) the neutron intruder states 1h11/2 instead of 1i13/2 and (ii) the proton spin-orbit partner states 1g7/2 rather than the intruder states 1h11/2. The goal is to identify and compare the principal features of nucleon polarisation effects for parallel and antiparallel coupling of spin and orbital angular momenta, respectively, and to investigate the properties of the selected orbitals upon crossing the magic neutron shell $N=82$.
@@
Non-perturbative shape and angular-momentum polarisations are the main features that distinguish the nuclear DFT from approaches based on valence-space approximations. Within the DFT picture, the unpaired nucleon determines the properties of an odd nucleus in the following way. For small quadrupole deformations near the semi-magic systems, an $I>1/2$ hole (particle) in the magnetic sub-state having the largest angular-momentum projection $\Omega$ equal to the total angular momentum $I$, $\Omega=I$, induces the prolate (oblate) polarisation of the nucleus. This rule is an immediate consequence of the geometrically driven dependence of the orbital single-particle energy $\epsilon$ on the axial deformation $\beta$, that is, $\epsilon\sim\beta\left(3\Omega^{2}-I(I+1)\right)$. Indeed, the orbital $\Omega=I$ then has the highest (lowest) energy $\epsilon\sim\beta{I}(2I-1)$ for $\beta>0$ ($\beta<0$) and thus determines the shape of the hole (particle) state.
@@
Analogous to the self-consistent shape polarisations, the non-zero magnetic dipole moment of the odd nucleon induces the non-zero magnetic dipole moments on the core nucleons, which, in turn, influences the magnetic dipole moment of the odd nucleon. In the self-consistent solution, these mutual angular-momentum polarisations (spin and orbital alike) produce the total magnetic dipole moment $\boldsymbol{\mu}$ of the odd nucleus in the same direction as the magnetic dipole moment of the odd nucleon. The time-odd mean-field sector of the energy functional determines the magnetic polarisation strength of the odd hole or particle. If the time-odd mean-field sector is neglected, only the magnetic moment of the odd nucleon survives. Again, since the mutual polarisations act throughout the full space of single-particle states, no effective $g$-factors are needed. As shown in [6], this feature of nuclear DFT removes the adjustable effective parameters customarily used in valence-space approaches.
For axial shapes, and especially for determining the magnetic moments, the nuclear DFT’s polarisation picture requires the following two elements:
@@
In [5, 6], we presented a brief literature overview of theoretical approaches used to determine nuclear electromagnetic moments; see also a recent review article [1]. Since then, the work in this direction has continued in several publications, such as the shell model [7][12], self-consistent Theory of Finite Fermi Systems (TFFS) [13, 14], Quasiparticle Phonon Nuclear Model (QPNM) [15], interacting boson-fermion model with configuration mixing [16, 17], Hartree-Fock-Bogolyubov (HFB) [18, 19, 20] and Hartree-Fock plus BCS (HF+BCS) approaches [21], ab initio no-core shell-model [22, 23], ab initio valence-space in-medium similarity renormalisation group (VS-IMSRG) [9, 11, 24], covariant density-functional theory (CDFT) [8], generator coordinate method within the multireference covariant density-functional theory (MR-CDFT) [25], multi-reference energy density functional (MR-EDF) calculations [26], and rigid triaxial rotor plus particle model [27]. In parallel, in [28][33], the methodology used in the present study was tested against the experimental data.
In [9][18], the results were obtained by employing effective charges and/or $g$-factors, with disparate values and justifications. Numerous diverse physical effects could be hidden and mixed under such numerical constants. These could include a limited size of the single-particle phase space, inadequate shape and/or angular-momentum polarisations, restrictions on conserved/broken symmetries or shape-alignment orientations, missing terms or deficiencies of the interactions or functionals, triaxial and/or octupole deformability, contributions from two-body meson-exchange currents, or maybe even a few more unknown unknowns, which potentially can all influence deviations between theory and experiment and deserve dedicated studies.
The paper is organised as follows. In section 2, we present details of the computational scheme employed in this work, and in section 3, we present the results obtained. Three subsections of section 4 discuss various aspects of the accuracy of the calculations related to the harmonic oscillator (HO) basis used, section 4.1, pairing strengths, section 4.2, and effects of particle number projection (PNP), section 4.3. Conclusions and outlook are given in section 5.
@@ 2 Methodology
    In the next step, we broke the spherical and time-reversal symmetries of ${}^{100}\text{Sn}$ by introducing small axial quadrupole constraints $\bar{Q}_{20}$ of about ±$\pm 0.9$ b, and a constraint on the $z$-component of the angular angular momentum ${J_{z}}$ for a small cranking frequency $\omega_{z}=0.001$  $\hbar$. The two quadrupole constraints gave us weakly deformed prolate and oblate ${}^{100}\text{Sn}$ states with the projections of the single-particle angular momenta quantised along the axial-symmetry $z$-axis. In this way, in the spectrum of single-particle states, we could identify two Nilsson orbitals, $\nu[505]11/2^{-}$ and $\pi[404]7/2^{+}$, which originate from the spherical neutron 1h11/2 and proton 1g7/2 states, respectively. The prolate and oblate single-particle states defined in this way were fixed and used to tag blocked quasiparticles for all open-shell odd nuclei considered in this work. In what follows, we refer to them as prolate and oblate tags.
@@
    The tagging mechanism is based on the strong affinity between the self-consistent quasiparticle states and single-particle states near closed shells. This affinity is the crux of our method for following specific configurations along many particle numbers.
@@
    In the next step, for each considered nucleus, depicted in figure 1, we performed tagged quasiparticle-blocking calculation by setting the average constant neutron and proton pairing gaps to 1 MeV [36]. We also set the axial quadrupole constraints for prolate and oblate tags to large values of $\bar{Q}_{20}=+10$ and $-10$ b, respectively. This allowed us to determine stable starting reference solutions suitable for obtaining the final self-consistent states.
@@
    In the next step, we determined the final self-consistent states by substituting the fixed pairing gaps with the neutron and proton pairing forces and relaxing the quadrupole-moment constraints. In this step, the pairing correlations were self-consistently determined. At the same time, during the self-consistent iteration, for prolate (oblate) tags, the quadrupole deformations were sliding back to lower quadrupole deformations down to prolate (oblate) minima. Both tags led to the same self-consistent solution in all cases where only a single minimum existed, whether prolate or oblate.
@@
    In the last step, for all self-consistent solutions, we restored the rotational symmetry [37] by projecting the intrinsic states on good angular momenta of $I=11/2$ or $7/2\,\hbar$ for the $\nu 11/2^{-}$ and $\pi 7/2^{+}$ configurations, respectively. At this point, the preserved axial symmetry and fixed quantised values of the angular momentum projections along the axial symmetry axis allowed for reducing the 3D integration over the three Euler angles [37] to the 1D integration over the rotation about the axis perpendicular to the axial symmetry axis only. This reduces the CPU time required to perform the angular momentum projection (AMP) by about two orders of magnitude and makes it comparable to that required to converge the HFB states. The spectroscopic electric quadrupole and magnetic dipole moments were then determined for symmetry-restored states.
The methodology proposed for the first time in [5] and outlined in this section allows us to perform large-scale calculations of nuclear electromagnetic moments for many elements and long chains of isotopes, from near-closed shells to open-shell configurations. The tagging mechanism allows for following given microscopic structures and configurations irrespective of their precise excitation energies and ordering in the spectra of odd nuclei. In this way, particular systematic agreement/disagreement patterns with data can be identified, and improvements in theoretical modelling can be proposed, implemented, studied, and tested.
@@ 3 Results
In this work, we employ the UNEDF1 [38] parametrisation of the Skyrme functional and focus on the time-odd mean fields generated by the spin-spin interactions of the form $\boldsymbol{\sigma}_{1}\cdot\boldsymbol{\sigma}_{2}$. The coupling constants of these interactions were parameterised by the standard isoscalar and isovector Landau parameters, $g_{0}$ and $g^{\prime}_{0}$ [39]. Following the analysis performed in [6], we used the values of these parameters fixed at $g_{0}=0.4$ and $g^{\prime}_{0}=1.7$. For the functional UNEDF1, in the notation defined in [40, 41], these values of the Landau parameters correspond to the Skyrme-functional coupling constants of $C_{0}^{s}=30.555$ and $C_{1}^{s}=129.858$ MeV fm3. In addition, the coupling constants that define the time-odd mean fields were either kept fixed at their values given by the gauge-symmetry conditions [40, 41] or set to zero, $C_{0}^{\Delta{s}}=C_{1}^{\Delta{s}}=0$ to avoid known time-odd instabilities [42]. In the pairing channel, we used the mixed pairing force adjusted in [38], with the strength parameters increased by 20% to account for replacing the original Lipkin-Nogami method with the HFB approach.
@@
As it turned out, almost all calculations performed for both tags (prolate and oblate) and for all 160 odd-$N$ and 147 odd-$Z$ nuclei (figure 1) converged well. Isotopes for which the oblate tag calculations did not converge were 115Ba and 157Sm. Similarly, the prolate tag calculations for 115I, 145I, 127Cs and 157Pm did not converge either.
@@
Figure 4: Energy differences between oblate tag and prolate tag minima for odd-$N$ (a) and odd-$Z$ (b) nuclei. The energy difference $E_{\text{oblate-tag}}-E_{\text{prolate-tag}}$ for 145I isotope ($N=92$) was linearly interpolated using the $E_{\text{prolate-tag}}$ of 143I and 147I with $N=90$ and $N=94$, respectively, the latter being outside the set of isotopes defined in figure 1.
In figure 4, for odd-$N$ (a) and odd-$Z$ (b) solutions, we show the energy differences between the oblate tag and prolate tag results. Vanishing values of those differences indicate cases where the prolate and oblate tags led to the unique (prolate or oblate) global minima. Results of that kind were obtained for (i) all Sn isotopes, (ii) all Te isotopes apart from $57\leq{N}\leq 63$, and all odd-$N$ elements with $77\leq{N}\leq 87$. Similarly, they were also obtained for (i) all Sb isotopes, (ii) all I isotopes apart from $50\leq{N}\leq 62$ and $N=92$, and (iii) all odd-$Z$ elements with $78\leq{N}\leq 86$. Other odd nuclei exhibit two different minima with positive (negative) oblate-prolate energy differences, which implies prolate (oblate) global minima.
Figure 5: Same as in figure 2 but for the global minima.
Figure 6: Same as in figure 3 but for the global minima.
It should be noted that this analysis is limited to axial shapes. Consequently, for small oblate-prolate energy differences, triaxial shapes may be relevant, which is beyond the scope of the current study. In the following, we present results solely for the global minima (unique, prolate tag, or oblate tag), and the Tables in the Supplemental Material [45] provide information on the type of minimum selected. The spectroscopic quadrupole moments at the global minima of odd-$N$ and odd-$Z$ nuclei are illustrated in figures 5 and 6, respectively.
@@
Figure 10: Calculated spectroscopic quadrupole moments compared with the intrinsic moments for the $\nu 11/2^{-}$ configurations in odd-$N$ nuclei (a) and $\pi 7/2^{+}$ configurations in odd-$Z$ (b) nuclei.
@@
We investigated the importance of the AMP in determining the spectroscopic electric quadrupole moments by testing the validity of the large-axial-deformation approximation [^1] See [35], Eq. (11.143)., which for $\Omega=I$ relates the spectroscopic and intrinsic multipole moments as,
$Q^{\text{spec}}_{\lambda}\simeq\left(C^{II}_{II,\lambda 0}\right)^{2}\times Q^{\text{intr}}_{\lambda 0},$ (1)
where $C^{II}_{II,\lambda 0}$ are the Clebsch-Gordan coefficients. For quadrupole moments ($Q=Q^{\text{spec}}_{2}$) and angular momenta $I=11/2$ and $7/2\,\hbar$, the corresponding Clebsch-Gordan coefficients have the following values,
$C^{\tfrac{11}{2}\tfrac{11}{2}}_{\tfrac{11}{2}\tfrac{11}{2},20}=\sqrt{\frac{55}{91}},\;\;\;\text{and}\;\;\;C^{\tfrac{7}{2}\tfrac{7}{2}}_{\tfrac{7}{2}\tfrac{7}{2},20}=\sqrt{\frac{7}{15}}.$ (2)
Figure 11: Same as in figure 10 but for the relative electric quadrupole moments (1).
@@
In panels (a) and (b) of figure 10, for the $\nu 11/2^{-}$ and $\pi 7/2^{+}$ configurations in odd-$N$ and odd-$Z$ nuclei, respectively, we compare the calculated spectroscopic electric quadrupole moments with their large-axial-deformation approximations (1). We see that the approximations seem to be perfect in the large scale of the figure. However, to visualise their precision at small deformations, in figure 11, we show the corresponding relative percentage deviations, $\left((C^{II}_{II,20})^{2}\times Q^{\text{intr}}_{20}-Q^{\text{spec}}_{2}\right)/Q^{\text{spec}}_{2}$. Here we see that for the $\nu 11/2^{-}$ configurations in odd-$N$ nuclei, at all deformations, the large-axial-deformation approximation is precise up to about 1%, whereas for the $\pi 7/2^{+}$ configurations in odd-$Z$ nuclei it deteriorates at small deformations to about 35%.
@@
Figure 12: Comparison between the calculated spectroscopic magnetic dipole moments $\mu$ of the $\nu 11/2^{-}$ configurations in odd-$N$ nuclei with experimental data [46, 54, 49, 55]. The uncertainties of the theoretical values were estimated based on the uncertainty of the Landau parameter $g^{\prime}_{0}=1.7(4)$ estimated in [6]. The asterisks () represent the cases where the signs of the magnetic dipole moments were not determined experimentally, and the calculated signs were adopted.

### v2

# Electromagnetic moments in the Sn-Gd region determined within nuclear DFT

@@ Abstract
Within the nuclear DFT framework, employing the Skyrme UNEDF1 functional and incorporating pairing correlations, we determined the spectroscopic electric quadrupole and magnetic dipole moments of the $\nu 11/2^{-}$ and $\pi 7/2^{+}$ configurations in heavy, deformed, open-shell odd nuclei with $50\leq Z\leq 64$. The notions of self-consistent shape and spin polarisations due to odd nucleons responsible for generating total electric quadrupole and magnetic dipole moments were transformed into detailed computational procedures. The alignment of intrinsic angular momentum along the axial symmetry axis, necessitating signature and time-reversal symmetry breaking, followed by the restoration of rotational symmetry, proved to be essential components of the method. In contrast, the restoration of particle number symmetry yields modifications of only about 1%. With the isovector spin-spin terms of the functional previously adjusted in near doubly magic nuclei across the mass chart, the calculations were parameter-free. Effective charges and $g$-factors were not employed. A reasonably good agreement was achieved between the calculated and measured electric quadrupole moments. A similarly fair description of the magnetic dipole moments was obtained for the intruder configurations $\nu 11/2^{-}$ alongside a poor description of those for $\pi 7/2^{+}$.
@@ 1 Introduction
In the present work, we expand the scope of the results presented in [5] to encompass the region of lighter elements, ranging from tin to gadolinium. Within nuclear density functional theory (DFT), we determine the spectroscopic electric quadrupole and magnetic dipole moments of the nuclei depicted in figure 1. Our dual focus here is on studying (i) the neutron intruder states 1h11/2 instead of 1i13/2 and (ii) the proton spin-orbit partner states 1g7/2 rather than the intruder states 1h11/2. The goal is to identify and compare the principal features of nucleon polarisation effects for parallel and antiparallel coupling of spin and orbital angular momenta, respectively, and to investigate the properties of the selected orbitals upon crossing the magic neutron shell $N=82$. The abundance of experimental data, e.g., the new experimental data on the electric quadrupole and magnetic dipole moments in tin isotopes [6] and antimony isotopes [7], serves as an advantage of studying the neutron 1h11/2 and proton 1g7/2 configurations.
@@
Non-perturbative shape and angular-momentum polarisations are the main features that distinguish the nuclear DFT from approaches based on valence-space approximations. Within the DFT picture, the unpaired nucleon determines the properties of an odd nucleus in the following way. For small quadrupole deformations near the semi-magic systems, an $I>1/2$ hole (particle) in the magnetic sub-state having the largest angular-momentum projection $\Omega$ equal to the total angular momentum $I$, $\Omega=I$, induces the prolate (oblate) polarisation of the nucleus. This rule is an immediate consequence of the geometrically driven dependence of the orbital single-particle energy $\epsilon$ on the axial deformation $\beta$, that is, $\epsilon\sim\beta\left(3\Omega^{2}-I(I+1)\right)$.
Such dependence on $I$ and $\Omega$ originates from the expectation value of a small quadrupole perturbation, $\hat{H}_{1}\sim\beta\hat{Q}_{20}$, of the standard spherical mean-field eigenstates $|N\ell{}I\Omega\rangle$. From the Wigner-Eckart theorem, the expectation value $\langle{}N\ell{}I\Omega|\hat{Q}_{20}|N\ell{}I\Omega\rangle$ of the quadrupole perturbation is proportional to the Clebsch-Gordan coefficient $C^{I\Omega}_{I\Omega,20}=\left(3\Omega^{2}-I(I+1)\right)/\left((2I-1)I(I+1)(2I+3)\right)$ [8]. Indeed, the orbital $\Omega=I$ then has the highest (lowest) energy $\epsilon\sim\beta{I}(2I-1)$ for $\beta>0$ ($\beta<0$) and thus determines the shape of the hole (particle) state.
@@
Analogous to the self-consistent shape polarisations, the non-zero magnetic dipole moment of the odd nucleon induces non-zero magnetic dipole moments in the core nucleons, which, in turn, influence the magnetic dipole moment of the odd nucleon. In the self-consistent solution, these mutual angular-momentum polarisations (spin and orbital alike) produce the total magnetic dipole moment $\boldsymbol{\mu}$ of the odd nucleus in the same direction as the magnetic dipole moment of the odd nucleon. The time-odd mean-field sector of the energy functional determines the magnetic polarisation strength of the odd hole or particle. If the time-odd mean-field sector is neglected, only the magnetic moment of the odd nucleon survives. Again, since the mutual polarisations act throughout the full space of single-particle states, no effective $g$-factors are needed. As shown in [9], this feature of nuclear DFT removes the adjustable effective parameters customarily used in valence-space approaches.
For axial shapes, and especially for determining the magnetic dipole moments $\boldsymbol{\mu}$, the nuclear DFT’s polarisation picture requires the following two elements:
@@
In [9, 5], we presented a brief literature overview of theoretical approaches used to determine nuclear electromagnetic moments; see also a recent review article [1]. Since then, the work in this direction has continued in several publications, such as the shell model [10][14], self-consistent Theory of Finite Fermi Systems (TFFS) [15, 16], Quasiparticle Phonon Nuclear Model (QPNM) [17], interacting boson-fermion model with configuration mixing [18, 19], Hartree-Fock-Bogolyubov (HFB) [20, 21, 22] and Hartree-Fock plus BCS (HF+BCS) approaches [23], ab initio IMSRG with projected generator coordinate method (PGCM) [24], ab initio in-medium no-core shell model (IM-NSCM) with PGCM [25], ab initio no-core shell-model [26, 27], ab initio valence-space in-medium similarity renormalisation group (VS-IMSRG) [12, 7, 28], covariant density-functional theory (CDFT) [11], generator coordinate method within the multireference covariant density-functional theory (MR-CDFT) [29], multi-reference energy density functional (MR-EDF) calculations [30], and rigid triaxial rotor plus particle model [31]. In parallel, the methodology used in the present study was tested against experimental data in [32][37].
In [12][20], the results were obtained by employing effective charges and/or $g$-factors, with disparate values and justifications. Numerous diverse physical effects could be hidden and mixed under such numerical constants. These could include a limited size of the single-particle phase space, inadequate shape and/or angular-momentum polarisations, restrictions on conserved or broken symmetries or shape-alignment orientations, missing terms or deficiencies in the interactions or functionals, triaxial and/or octupole deformability, contributions from two-body meson-exchange currents, or maybe even a few more unknown unknowns, which potentially can all influence deviations between theory and experiment and deserve dedicated studies.
The paper is organised as follows. In section 2, we present details of the computational scheme employed in this work, and in section 3, we present the results obtained. Three subsections of section 4 discuss various aspects of the precision of the calculations related to the harmonic oscillator (HO) basis used, section 4.1, pairing strengths, section 4.2, and effects of particle number projection (PNP), section 4.3. Conclusions and outlook are presented in section 5.
@@ 2 Methodology
    In the next step, we broke the spherical and time-reversal symmetries of ${}^{100}\text{Sn}$ by introducing small axial quadrupole constraints $\bar{Q}_{20}$ of about ±$\pm 0.9$ b, and a constraint on the $z$-component of the angular angular momentum ${J_{z}}$ for a small cranking frequency $\omega_{z}=0.001$  $\hbar$. The constraint on $J_{z}$ (using a small cranking frequency $\omega_{z}$) enforces the alignment of the angular momenta along the axis of axial symmetry of the deformed nucleus. Together with the constraint on $J_{z}$, the two quadrupole constraints gave us weakly deformed prolate and oblate ${}^{100}\text{Sn}$ states with the projections of the single-particle angular momenta quantised along the axial-symmetry $z$-axis. In this way, in the spectrum of single-particle states, we could identify two Nilsson orbitals, $\nu[505]11/2^{-}$ and $\pi[404]7/2^{+}$, which originate from the spherical neutron 1h11/2 and proton 1g7/2 states, respectively. The prolate and oblate single-particle states defined in this way were fixed and used to tag blocked quasiparticles for all open-shell odd nuclei considered in this work. In what follows, we refer to them as prolate and oblate tags.
@@
    The tagging mechanism is based on the strong affinity between the self-consistent quasiparticle states and single-particle states near closed shells. This affinity is the crux of our method for following specific configurations across a range of particle numbers.
    In summary, we first defined the fixed single-particle states $\nu[505]11/2^{-}$ and $\pi[404]7/2^{+}$ by applying the quadrupole constraints $\bar{Q}_{20}=\pm 0.9$ b. Then, we selected the one-quasiparticle state by maximising the overlap $\mathcal{O}_{\mu}$ between the quasiparticle wave functions and the previously defined fixed single-particle states $\nu[505]11/2^{-}$ and $\pi[404]7/2^{+}$.
@@
    In the next step, for each considered nucleus, depicted in figure 1, we performed a tagged quasiparticle-blocking calculation by setting the average constant neutron and proton pairing gaps to 1 MeV [40]. We also imposed axial quadrupole constraints for prolate and oblate tags with large values of $\bar{Q}_{20}=+10$ and $-10$ b, respectively. This allowed us to determine stable starting reference solutions suitable for obtaining the final self-consistent states.
@@
    In the next step, we determined the final self-consistent states by substituting the fixed pairing gaps with the neutron and proton pairing forces and relaxing the quadrupole-moment constraints. In this step, the pairing correlations were self-consistently determined. At the same time, during the self-consistent iteration, for both prolate and oblate tags, the quadrupole deformations slid back to lower quadrupole deformations, reaching the prolate and oblate minima, respectively. Both tags led to the same self-consistent solution in all cases where only a single minimum existed, whether prolate or oblate.
@@
    In the last step, for all self-consistent solutions, we restored the rotational symmetry [41] by projecting the intrinsic states on good angular momenta of $I=11/2$ or $7/2\,\hbar$ for the $\nu 11/2^{-}$ and $\pi 7/2^{+}$ configurations, respectively. At this point, the preserved axial symmetry and fixed quantised values of the angular momentum projections along the axial symmetry axis allowed for reducing the 3D integration over the three Euler angles [41] to the 1D integration over the rotation about the axis perpendicular to the axial symmetry axis only. This reduces the CPU time required to perform the angular momentum projection (AMP) by about two orders of magnitude and makes it comparable to that needed to converge the HFB states. The spectroscopic electric quadrupole and magnetic dipole moments were then determined for symmetry-restored states.
The methodology proposed for the first time in [5] and outlined in this section enables us to perform large-scale calculations of nuclear electromagnetic moments for many elements and long chains of isotopes, ranging from near-closed shells to open-shell configurations. The tagging mechanism allows us to follow given microscopic structures and configurations irrespective of their precise excitation energies and ordering in the spectra of odd nuclei. In this way, particular systematic agreement or disagreement patterns with data can be identified, and improvements in theoretical modelling can be proposed, implemented, studied, and tested.
@@ 3 Results
In this work, we employ the UNEDF1 [42] parametrisation of the Skyrme functional and focus on the time-odd mean fields generated by the spin-spin interactions of the form $\boldsymbol{\sigma}_{1}\cdot\boldsymbol{\sigma}_{2}$. The standard Skyrme functionals, like UNEDF1, generate the time-odd mean fields via the following terms [43, 44]: $C^{2}_{t}s^{2}_{t}(\textbf{r})$, $C^{\Delta s}_{t}\textbf{s}_{t}(\textbf{r})\cdot\nabla^{2}\textbf{s}_{t}(\textbf{r})$, $C^{T}_{t}\textbf{s}_{t}(\textbf{r})\cdot\textbf{T}_{t}(\textbf{r})$, $C^{j}_{t}j^{2}_{t}(\textbf{r})$, and $C^{\nabla j}_{t}\textbf{s}_{t}(\textbf{r})\cdot[\boldsymbol{\nabla}\times\textbf{j}_{t}](\textbf{r})$, where $t=0(1)$ refers to the isoscalar (isovector) character of the coupling constants, $\textbf{s}_{t}(\textbf{r})$ is the spin density, $\textbf{T}_{t}(\textbf{r})$ represents the vector kinetic density, and $\textbf{j}_{t}(\textbf{r})$ represents the current density. As in our previous work [9, 5], we parameterised coupling constants $C^{s}_{t}$ and $C^{T}_{t}$ by the standard isoscalar and isovector Landau parameters, $g_{0}$ and and $g^{\prime}_{0}$ [45], and we used the values of $g_{0}=0.4$ and $g^{\prime}_{0}=1.7$. For the functional UNEDF1, these values of the Landau parameters correspond to the Skyrme-functional coupling constants of $C^{s}_{0}=30.555$ and $C^{s}_{1}=129.858$ MeV fm3. In addition, the coupling constants $C^{T}_{t}$, $C^{j}_{t}$, and $C^{\nabla j}_{t}$ were kept fixed at their values given by the gauge symmetry conditions [43, 44] and $C^{\Delta s}_{t}$ were set to zero to avoid known time-odd instabilities [46].
As in our previous work [5], in the pairing channel, we used the mixed pairing force adjusted in [42], and we increased the strength parameters by 20% to account for replacing the original Lipkin-Nogami (LN) method with the HFB approach. Within the LN method and the BCS model, see, e.g., [47], the approximate particle-number projected spectral gaps are given by $\langle v^{2}\Delta\rangle^{(\text{LN})}_{q}=\langle v^{2}\Delta\rangle_{q}+\lambda_{2,q}$ and $\langle uv\Delta\rangle^{(\text{LN})}_{q}=\langle uv\Delta\rangle_{q}+\lambda_{2,q}$, where $q\in\{p,n\}$, and are larger by the LN second-order coefficient $\lambda_{2,q}$ than the corresponding BCS gaps. Although the discussion in [47] centres on pairing gaps within the BCS approximation, a similar conclusion holds for the HFB approach used here because, within a canonical basis [39], the BCS approximation is equivalent to the HFB approach. Since the calculated pairing gaps increase with the pairing strengths, the LN contributions of $\lambda_{2}$ to the spectral gaps can be compensated by increasing the HFB pairing strengths. A 20% increase was chosen so that our calculated HFB values of the odd-even mass staggering (section 4.2) are, on average, compatible with experimental data.
@@
As it turned out, almost all calculations performed for both tags (prolate and oblate) and for all 160 odd-$N$ and 147 odd-$Z$ nuclei (figure 1) converged well. Isotopes for which the oblate-tag calculations did not converge were 115Ba and 157Sm. Similarly, the prolate-tag calculations for 115I, 145I, 127Cs and 157Pm did not converge either.
@@
Figure 4: Energy differences between oblate-tag and prolate-tag minima for odd-$N$ (a) and odd-$Z$ (b) nuclei. The energy difference $E_{\text{oblate-tag}}-E_{\text{prolate-tag}}$ for 145I isotope ($N=92$) was linearly interpolated using the $E_{\text{prolate-tag}}$ of 143I and 147I with $N=90$ and $N=94$, respectively, the latter being outside the set of isotopes defined in figure 1.
In figure 4, for odd-$N$ (a) and odd-$Z$ (b) solutions, we show the energy differences between the oblate-tag and prolate-tag results. Vanishing values of those differences indicate cases where the prolate and oblate tags led to the unique (prolate or oblate) global minima. Results of that kind were obtained for (i) all Sn isotopes, (ii) all Te isotopes apart from $57\leq{N}\leq 63$, and all odd-$N$ elements with $77\leq{N}\leq 87$. Similarly, they were also obtained for (i) all Sb isotopes, (ii) all I isotopes apart from $50\leq{N}\leq 62$ and $N=92$, and (iii) all odd-$Z$ elements with $78\leq{N}\leq 86$. Other odd nuclei exhibit two different minima with positive (negative) oblate-prolate energy differences, which implies prolate (oblate) global minima.
Figure 5: Same as in figure 2 but for the global minima (a) and the corresponding Bohr parameters $\beta_{2}$ (b) calculated using Eq. (1).
Figure 6: Same as in figure 3 but for the global minima (a) and the corresponding Bohr parameters $\beta_{2}$ (b) calculated using Eq. (1).
It should be noted that this analysis is limited to axial shapes. Consequently, for small oblate/prolate energy differences or small deformations, triaxial shapes and the generator coordinate mixing thereof leading to shape coexistence may be relevant, which is beyond the scope of the current study. In the following, we present results solely for the global minima (unique, prolate-tag, or oblate-tag), and the Tables in the Supplemental Material [50] provide information on the type of minimum selected.
In figures 5(a) and 6(a), we present the spectroscopic quadrupole moments $Q$ calculated at the global minima of odd-$N$ and odd-$Z$ nuclei, respectively. For illustration, in figures 5(b) and 6(b), we also present the corresponding Bohr $\beta_{2}$ parameters, calculated from the values of $Q$ in an analogous way as they are routinely determined from the measured spectroscopic quadrupole moments, that is,
$\beta_{2}=\frac{\sqrt{5\pi}}{3ZR^{2}_{0}}\frac{Q}{\left(C^{II}_{II,20}\right)^{2}},$ (1)
[…]
@@
Figure 10: Calculated spectroscopic quadrupole moments $Q$ compared with estimates (2) for the $\nu 11/2^{-}$ configurations in odd-$N$ nuclei (a) and $\pi 7/2^{+}$ configurations in odd-$Z$ (b) nuclei.
@@
We investigated the importance of the AMP in determining the spectroscopic electric quadrupole moments by testing the validity of the large-axial-deformation (or rigid-rotor) approximation [^1] See [39], Eq. (11.143)., already used in (1). It allows for estimating the spectroscopic quadrupole moments $Q^{\text{spec}}_{\text{est}}$ from the calculated intrinsic quadrupole moments $Q^{\text{intr}}_{20}$ as,
$Q^{\text{spec}}_{\text{est}}=\left(C^{II}_{II,20}\right)^{2}\times Q^{\text{intr}}_{20}.$ (2)
Figure 11: Same as in figure 10 but for the relative electric quadrupole moments, see text.
@@
In panels (a) and (b) of figure 10, for the $\nu 11/2^{-}$ and $\pi 7/2^{+}$ configurations in odd-$N$ and odd-$Z$ nuclei, respectively, we compare the calculated spectroscopic electric quadrupole moments $Q$ with their rigid-rotor approximations (2). We observe that, when plotted in the large scale used in the figure, the approximations seem to be perfect. However, to visualise their precision at small deformations, in figure 11, we show the corresponding relative percentage deviations, $\left(Q^{\text{spec}}_{\text{est}}-Q\right)/Q$. Here we see that for the $\nu 11/2^{-}$ configurations in odd-$N$ nuclei, at all deformations, the rigid-rotor approximation is precise up to about 1%, whereas for the $\pi 7/2^{+}$ configurations in odd-$Z$ nuclei it deteriorates at small deformations to about 35%. Figure 11 may serve as a guide for employing the rigid-rotor approximation rather than conducting the exact AMP in weakly deformed nuclei.
@@
Figure 12: Comparison between the calculated spectroscopic magnetic dipole moments $\mu$ of the $\nu 11/2^{-}$ configurations in odd-$N$ nuclei with experimental data [51, 58, 6, 59]. The uncertainties of the theoretical values were obtained by recalculating the theoretical results for the Landau parameters $g^{\prime}_{0}=1.3$ and $2.1$, which correspond to the uncertainty of $g^{\prime}_{0}=1.7(4)$ estimated in [9]. The asterisks () represent the cases where the signs of the magnetic dipole moments were not determined experimentally, and the calculated signs were adopted.
Figure 13: Same as in figure 12 but for the $\pi 7/2^{+}$ configurations in odd-$Z$ nuclei  [7].